# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 4.2. Deformation of special divisors on boundary charts
with exact rows and $H_{1}^{\mathrm{dR}}(M)$ is the pushout of $H_{1}^{\mathrm{dR}}(G^{\sharp})$ along $\mathrm{Fil}(G^{\sharp})\rightarrow\mathrm{Fil}(M)$. Define
(4.10) $L_{M}=\bar{\epsilon}\operatorname{Lie}M.$
Again we can easily see that $L_{M}$ is an $\mathcal{O}_{\mathcal{C}_{\Phi}}$-module local direct summand of rank one. It is stable under $\mathcal{O}_{\boldsymbol{k}}$, which acts on $\operatorname{Lie}M/L_{M}$ and $L_{M}$ via $\varphi,\bar{\varphi}$ : $\mathcal{O}_{\boldsymbol{k}}\rightarrow\mathcal{O}_{\mathcal{C}_{\Phi}}$, respectively. Using the canonical isomorphism $\operatorname{Lie}G^{\sharp}\cong\operatorname{Lie}M$, we can identify $L_{G^{\sharp}}$ with $L_{M}$ over $\mathcal{C}_{\Phi}$.
We now revisit the setting at the beginning of this section with $S=\mathcal{C}_{\Phi}^{*}$ and $Z=\mathcal{Z}_{\Phi}^{*}(m)$. For any geometric point $z\in\mathcal{C}_{\Phi}$, by results in [${}^{+}$20, §3.6], there is an étale neighborhood $U$ of $z$ in $\mathcal{C}_{\Phi}^{*}$ such that $\mathcal{Z}_{\Phi}^{*}(m)|_{U}\rightarrow\mathcal{C}_{\Phi}^{*}|_{U}$ is a closed immersion of schemes on every connected component of its domain. Let $\mathcal{Z}$ be such a connected component. The universal morphism $x:\mathbf{A}_{0}|_{\mathcal{Z}}\rightarrow G^{\sharp}|_{\mathcal{Z}}$ induces a morphism of vector bundles over $\mathcal{Z}$.
[…]
which maps $\mathrm{Fil}(\mathbf{A}_{0}|_{\mathcal{Z}})$ to $\mathrm{Fil}(G^{\sharp}|_{\mathcal{Z}})$. By Grothendieck-Messing theory for $1$-motives, this morphism admits a canonical extension to a morphism between vector bundles over $\tilde{\mathcal{Z}}$,
$H_{1}^{\mathrm{dR}}(\mathbf{A}_{0}|_{\tilde{\mathcal{Z}}})\xrightarrow{\tilde{x}}H_{1}^{\mathrm{dR}}(G^{\sharp}|_{\tilde{\mathcal{Z}}}),$
which determines a homorphism
(4.11) $\mathrm{Fil}(\mathbf{A}_{0}|_{\tilde{\mathcal{Z}}})\xrightarrow{\tilde{x}}\operatorname{Lie}(G^{\sharp}|_{\tilde{\mathcal{Z}}}).$
@@
(4.12) $\operatorname{Fil}\left(\mathbf{A}_{0}\right)|_{\tilde{\mathcal{Z}}}\stackrel{{\scriptstyle\tilde{x}}}{{\rightarrow}}L_{G^{\sharp}}|_{\tilde{\mathcal{Z}}}.$
@@
By the signature condition of $\mathbf{A}_{0}$, we know that $\bar{\epsilon}H_{1}^{\mathrm{dR}}(\mathbf{A}_{0})\subset\mathrm{Fil}(\mathbf{A}_{0})$. Since both of them are locally free of rank $1$, we know they are equal. Hence (4.11) takes values in $\bar{\epsilon}\operatorname{Lie}G^{\sharp}|_{\tilde{\mathcal{Z}}}$ which is exactly $L_{G^{\sharp}}|_{\tilde{\mathcal{Z}}}$ by definition. This proves the first claim of the proposition. The second claim follows from Grothendieck-Messing theory for $1$-motives, see for example [MP11, Lemma 1.3.4.1]. ∎
4.3. Deformation of $\mathcal{Z}^{*}(m)$ in $\mathcal{S}^{*}$
We now revisit the setting at the beginning of this section. Recall that for any geometric point $z\in\mathcal{S}$, we choose an étale neighborhood $U$ of $z$ in $\mathcal{S}$ such that $\mathcal{Z}(m)|_{U}\rightarrow\mathcal{S}|_{U}$ is a closed immersion of schemes on every connected component of its domain. Let $\mathcal{Z}$ be such a connected component. Let $U^{*}$ be the Zariski closure of $U$ in $\mathcal{S}^{*}$, and $\mathcal{Z}^{*}$ be the Zariski closure of $\mathcal{Z}$ in $\mathcal{S}^{*}$. Recall that $\tilde{\mathcal{Z}}$ is the first order thickening of $\mathcal{Z}$ in $\mathcal{S}|_{U}$ as defined at the beginning of the section. Analogously let $\tilde{\mathcal{Z}}^{*}$ be the first order thickening of $\mathcal{Z}^{*}$ in $\mathcal{S}^{*}|_{U^{*}}$. The following is the main result of this section.
Theorem 4.4.
The morphism (4.2) can be extended to a morphism between line bundles over $\mathcal{S}^{*}$:
(4.13) $\operatorname{Fil}\left(\mathbf{A}_{0}\right)|_{\tilde{\mathcal{Z}}^{*}}\stackrel{{\scriptstyle\tilde{x}}}{{\rightarrow}}L_{\mathbf{G}}|_{\tilde{\mathcal{Z}}^{*}}.$
The Kudla-Rapoport divisor $\mathcal{Z}^{*}$ is the largest closed formal subscheme of $\tilde{\mathcal{Z}}^{*}$ over which (4.13) is trivial.
@@ 4.3. Deformation of $\mathcal{Z}^{*}(m)$ in $\mathcal{S}^{*}$
Without loss of generality, by shrinking $U$ if necessary we can assume that $U^{*}$ intersects with exactly one boundary component $\mathcal{B}_{\Phi}$ nontrivially. As before assume assume $\Phi=(\mathfrak{n},L_{\Phi})$ and $\mathfrak{m}$ be the hermitian dual of $\mathfrak{n}$. Recall from Theorem 2.5 that we have a morphism of stacks
$\pi_{\Phi}:(\mathcal{C}_{\Phi}^{*})^{\wedge}_{\mathcal{B}_{\Phi}}\rightarrow(\mathcal{S}^{*})^{\wedge}_{\mathcal{S}^{*}(\Phi)},$
which is a $\Delta_{\Phi}$-torsor and maps $\mathcal{B}_{\Phi}$ to $\mathcal{S}^{*}(\Phi)$. By Proposition 4.1, it suffices to assume that $U^{*}$ is an étale neighborhood of $(\mathcal{S}^{*})^{\wedge}_{\mathcal{S}^{*}(\Phi)}$ and $U=U^{*}\setminus\mathcal{S}^{*}(\Phi)|_{U^{*}}$. Let $\mathcal{C}_{\Phi}^{\wedge}=(\mathcal{C}_{\Phi}^{*})^{\wedge}_{\mathcal{B}_{\Phi}}\setminus\mathcal{B}_{\Phi}$.
As explained in §4.1, we can associate a semi-abelian scheme $G^{\sharp}$ with $\mathcal{O}_{\boldsymbol{k}}$-action over $(\mathcal{C}_{\Phi}^{*})^{\wedge}_{\mathcal{B}_{\Phi}}$, and a principally polarized $1$-motive $M$ with $\mathcal{O}_{\boldsymbol{k}}$-action over $\mathcal{C}_{\Phi}^{\wedge}$. The relation between $G^{\sharp}$, $M$ and the pullback of $\mathbf{G}$ over $\mathcal{S}^{*}$ under $\pi_{\Phi}$ is as follows. Recall that we have an exact sequence of fppf sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules over $(\mathcal{C}_{\Phi}^{*})^{\wedge}_{\mathcal{B}_{\Phi}}$
$0\rightarrow\mathfrak{m}\otimes_{\mathbb{Z}}\mathbb{G}_{m}\rightarrow G^{\sharp}\rightarrow B\rightarrow 0,$
@@
By construction the polarizations of $A$ and $M$ are compatible, hence $\mathcal{Z}(m)$ pulls back to $\mathcal{Z}_{\Phi}(m)$ under the morphism $\pi_{\Phi}$, see [${}^{+}$20, Theorem 3.7.1(8)]. Taking Zariski closure, we see that $\mathcal{Z}^{*}(m)$ pulls back to $\mathcal{Z}_{\Phi}^{*}(m)$. Suppose that $\mathcal{Z}^{*}$ pulls back under $\pi_{\Phi}$ to $\mathcal{Y}^{*}$ on $(\mathcal{C}_{\Phi}^{*})^{\wedge}_{\mathcal{B}_{\Phi}}|_{U^{*}}$. Denote by $\mathcal{Y}$ the restriction of $\mathcal{Y}^{*}$ to $\mathcal{C}_{\Phi}^{\wedge}|_{U}$ and let their first order thickening in $(\mathcal{C}_{\Phi}^{*})^{\wedge}_{\mathcal{B}_{\Phi}}|_{U^{*}}$ be $\tilde{\mathcal{Y}}$ and $\tilde{\mathcal{Y}^{*}}$ respectively. By Lemma 4.5 below, we have a canonical isomorphism preserving weight filtration
$H_{1}^{\mathrm{dR}}(M)\cong H_{1}^{\mathrm{dR}}(A).$
Hence we can identify $\operatorname{Lie}A$ with $\operatorname{Lie}M$, $L_{A}$ with $L_{M}$ which is further identified with $L_{G^{\sharp}}$. Under these identifications, the obstruction morphism in (4.2)
$\operatorname{Fil}\left(\mathbf{A}_{0}\right)|_{\tilde{\mathcal{Z}}}\stackrel{{\scriptstyle\tilde{x}}}{{\rightarrow}}L_{\mathbf{A}}|_{\tilde{\mathcal{Z}}}$
pull backs under $\pi_{\Phi}$ to a morphism
$\operatorname{Fil}\left(\mathbf{A}_{0}\right)|_{\tilde{\mathcal{Y}}}\stackrel{{\scriptstyle\tilde{y}}}{{\rightarrow}}L_{M}|_{\tilde{\mathcal{Y}}}\cong L_{G^{\sharp}}|_{\tilde{\mathcal{Y}}}.$
By Proposition 4.3, this obstruction homomorphism extends to $\tilde{\mathcal{Y}}^{*}$, and $\mathcal{Y}^{*}$ is the vanishing locus of $\tilde{y}$ in $\tilde{\mathcal{Y}}^{*}$. The extended homomorphism is obviously equi-variant under the action of $\Delta_{\Phi}$, so descends to a homomorphism
$\operatorname{Fil}\left(\mathbf{A}_{0}\right)|_{\tilde{\mathcal{Z}}^{*}}\stackrel{{\scriptstyle\tilde{x}}}{{\rightarrow}}L_{\mathbf{G}}|_{\tilde{\mathcal{Z}}^{*}}.$
And $\mathcal{Z}^{*}$ is the vanishing locus of $\tilde{x}$ in $\tilde{\mathcal{Z}}^{*}$. This finishes the proof of the proposition. ∎
Lemma 4.5.
Let the notation be as above. Then we have a canonical identification of de Rham realizations
$H_{1}^{\mathrm{dR}}(M)\cong H_{1}^{\mathrm{dR}}(A)$
preserving weight filtrations.
@@
Let $E(A)$ (resp. $E(G^{\sharp})$) be the universal vector extension of $A$ (resp. $G^{\sharp}$). So we have an exact sequence of fppf-sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules
$0\rightarrow\mathrm{Fil}(A)=\underline{\mathrm{Ext}}^{1}(A,\mathbb{G}_{a})^{\vee}\rightarrow E(A)\rightarrow A\rightarrow 0$
Consider the fiber product $E:=E(A)\times_{A}G^{\sharp}$ where we use the natural quotient map $G^{\sharp}\rightarrow A$. Then $E$ is an extension of $G^{\sharp}$ by the vector group $\underline{\mathrm{Ext}}^{1}(A,\mathbb{G}_{a})^{\vee}$. Moreover the natural morphism $E\rightarrow E(A)$ is étale (as $G^{\sharp}\rightarrow A$ is) so induces an isomorphism on Lie algebras
$\operatorname{Lie}E\cong\operatorname{Lie}E(A)=H_{1}^{\mathrm{dR}}(A).$
By the universal property of $E(G^{\sharp})$, we have a commutative diagram of fppf-sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules
[…]
Taking Lie algebra, we have a commutative diagram of locally free sheaves
[…]
We would like to identify (4.14) with (4.9). In fact by the main result of [Ber05], we know that
$\mathrm{Fil}(A)\cong(\operatorname{Lie}A^{\vee})^{\vee}\cong(\operatorname{Lie}(G^{\sharp})^{\vee})^{\vee}=(\operatorname{Lie}M^{\vee})^{\vee}\cong\mathrm{Fil}(M).$
Moreover the maps $\mathrm{Fil}(G^{\sharp})\rightarrow\mathrm{Fil}(A)$ and $\mathrm{Fil}(G^{\sharp})\rightarrow\mathrm{Fil}(M)$ can be identified under this isomorphism. Hence $H_{1}^{\mathrm{dR}}(A)$ and $H_{1}^{\mathrm{dR}}(M)$ have to agree, both are pushout of $H_{1}^{\mathrm{dR}}(G^{\sharp})$ along $\mathrm{Fil}(G^{\sharp})\rightarrow\mathrm{Fil}(M)$. ∎
5. Pullback of algebraic divisors

### v2

@@ 4.2. Deformation of special divisors on formal boundary charts
We now revisit the setting at the beginning of this section with $S=\mathcal{C}_{\Phi}^{*}$ and $Z=\mathcal{Z}_{\Phi}^{*}(m)$. For any geometric point $z\in\mathcal{C}_{\Phi}$, there is an étale neighborhood $U$ of $z$ in $\mathcal{C}_{\Phi}^{*}$ such that $\mathcal{Z}_{\Phi}^{*}(m)|_{U}\rightarrow\mathcal{C}_{\Phi}^{*}|_{U}$ is a closed immersion of schemes on every connected component of its domain. Let $\mathcal{Z}$ be such a connected component. The universal morphism $x:A_{0}|_{\mathcal{Z}}\rightarrow G^{\sharp}|_{\mathcal{Z}}$ induces a morphism of vector bundles over $\mathcal{Z}$.
$H_{1}^{\mathrm{dR}}(A_{0}|_{\mathcal{Z}})\xrightarrow{x}H_{1}^{\mathrm{dR}}(G^{\sharp}|_{\mathcal{Z}}),$
which maps $\mathrm{Fil}(A_{0}|_{\mathcal{Z}})$ to $\mathrm{Fil}(G^{\sharp}|_{\mathcal{Z}})$. By Grothendieck-Messing theory for $1$-motives (see for example [AB11]), this morphism admits a canonical extension to a morphism between vector bundles over $\tilde{\mathcal{Z}}$,
@@
(4.9) $\operatorname{Fil}\left(A_{0}\right)|_{\tilde{\mathcal{Z}}}\stackrel{{\scriptstyle\tilde{x}}}{{\rightarrow}}L_{G^{\sharp}}|_{\tilde{\mathcal{Z}}}$.
@@
By the signature condition of $A_{0}$, we know that $\bar{\epsilon}H_{1}^{\mathrm{dR}}(A_{0})\subset\mathrm{Fil}(A_{0})$. Since both of them are rank 1 local direct summands of the vector bundle $H_{1}^{\mathrm{dR}}(A_{0})$, we know they are equal. Hence (4.8) takes values in $\bar{\epsilon}\operatorname{Lie}G^{\sharp}|_{\tilde{\mathcal{Z}}}$ which is exactly $L_{G^{\sharp}}|_{\tilde{\mathcal{Z}}}$ by definition. This proves the first claim of the proposition. The second claim follows from Grothendieck-Messing theory for $1$-motives, see for example [AB11, Theorem 2.1 (iv), Remark 2.2 (c)]. ∎
4.3. Special divisors on the toroidal compactification
The tautological degeneration data $(B,\kappa,\psi,c,c^{\vee})$ of type $\Phi$ and signature (n1,1)(n-1,1) over $\mathcal{S}^{*}(\Phi)$ determines a morphism of semi-abelian schemes $\lambda_{\mathbf{G}}:\mathbf{G}\rightarrow\mathbf{G}^{\vee}$ which extends the principal polarization $\lambda:\mathbf{A}\rightarrow\mathbf{A}^{\vee}$. Using $\lambda_{\mathbf{G}}$ we can define a hermitian form $(,)$ on $\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}\left(A_{0},G\right)$ in the same way as in (2.9). By Proposition 2.9 and Lemma 4.2, the form is positive definite.
Definition 4.5.
For any $m\in\mathbb{Z}_{>0}$, define $\mathcal{Z}^{*}(m)$ to be the stack assigning to each $\mathcal{O}_{\boldsymbol{k}}$-scheme $S$ with an $\mathcal{O}_{\boldsymbol{k}}$-morphism $S\rightarrow\mathcal{S}^{*}$ the set
$\mathcal{Z}^{*}(m)(S)=\{x\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}\left(A_{0},G\right)\mid(x,x)=m\},$
where $\left(A_{0},\iota_{0},\lambda_{0},G,\iota,\lambda,\mathcal{F}\right)$ is the pullback of the universal object over $\mathcal{S}^{*}$ to $S$.
Recall that we have the following rigidity lemma of semi-abelian schemes.
Lemma 4.6.
Let $S$ be a noetherian scheme and $S_{0}$ be its reduced locus. Assume $G$ and $H$ are semi-abelian schemes over $S$. Then the reduction map induces a canonical inclusion
$\operatorname{Hom}(G,H)\hookrightarrow\operatorname{Hom}(G|_{S_{0}},H|_{S_{0}}),$
where we take homomorphisms between group schemes.
@@ 4.3. Special divisors on the toroidal compactification
This follows easily from the rigidity lemma of abelian schemes and the fact that there are no non-constant homomorphisms from tori to abelian schemes. ∎
Proposition 4.7.
For each integer $m>0$, $\mathcal{Z}^{*}(m)$ is a Deligne-Mumford stack containing $\mathcal{Z}(m)$ as an open substack. The forgetful morphism $\mathcal{Z}^{*}(m)\rightarrow\mathcal{S}^{*}$ is finite and unramified.
Proof.
Given an $S$-valued point $\left(A_{0},\iota_{0},\lambda_{0},G,\iota,\lambda,\mathcal{F}\right)$ of $\mathcal{S}^{*}$, the functor $\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}\left(A_{0},G\right)$ on $(\mathrm{Sch}/S)$ defined by
$S^{\prime}\mapsto\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}\left(A_{0}\times_{S}S^{\prime},G\times_{S}S^{\prime}\right)$
is representable by a scheme by Grothendieck’s representability theorem on $\operatorname{Hom}$ sheaves. Given this, since $\mathcal{S}^{*}$ is a Deligne-Mumford stack, so is $\mathcal{Z}^{*}(m)$.
Unramifiedness of the forgetful morphism $\mathcal{Z}^{*}(m)\rightarrow\mathcal{S}^{*}$ follows from the infinitesimal criterion for unramifiedness and Lemma 4.6. Quasi-finiteness of the morphism $\mathcal{Z}^{*}(m)\rightarrow\mathcal{S}^{*}$ follows from the positivity of the hermitian form $(,)$. To prove the properness of the morphism $\mathcal{Z}^{*}(m)\rightarrow\mathcal{S}^{*}$, we use the valuation criterion. Assume that we have a commutative diagram of morphisms
$\mathrm{Spec}\,K$$\mathcal{Z}^{*}(m)$$\mathrm{Spec}\,V$$\mathcal{S}^{*}.$}
where $V$ is a DVR with fractional field $K$, we would like to show that there exists a dashed arrow $\mathrm{Spec}\,V\dasharrow\mathcal{Z}^{*}(m)$ after adding which the diagram still commutes. Let $\left(A_{0},\iota_{0},\lambda_{0},G,\iota,\lambda,\mathcal{F}\right)$ be the pullback of the universal object over $\mathcal{S}^{*}$ to $\mathrm{Spec}\,V$. By assumption there exists $x|_{K}\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(A_{0}|_{\mathrm{Spec}\,K},G|_{\mathrm{Spec}\,K})$ such that $(x|_{K},x|_{K})=m$. By [FC13, Chapter I, Proposition 2.7], $x|_{K}$ extends to a homomorphism $x\in\operatorname{Hom}(A_{0},G)$. As commuting with $\mathcal{O}_{\boldsymbol{k}}$-action is a closed condition, we have $x\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(A_{0},G)$. By similar reasoning, we have $(x,x)=m$. This shows that the dashed arrow does exist. ∎
By Proposition 4.7, for any geometric point $z\in\partial\mathcal{S}^{*}$, there is an étale neighborhood $U$ of $z$ in $\mathcal{S}^{*}$ such that $\mathcal{Z}^{*}(m)|_{U}\rightarrow\mathcal{S}^{*}|_{U}$ is a closed immersion of schemes on every connected component of its domain. Let $\hat{U}$ be the completion of $U$ along $\partial S^{*}$. Let $\mathcal{Z}$ be a connected component of $\mathcal{Z}^{*}(m)|_{\hat{U}}$ and $\tilde{\mathcal{Z}}$ be its first order thickening. The universal morphism $x:\mathbf{A}_{0}|_{\mathcal{Z}}\rightarrow\mathbf{G}|_{\mathcal{Z}}$ induces a morphism of vector bundles over $\mathcal{Z}$
$H_{1}^{\mathrm{dR}}(\mathbf{A}_{0}|_{\mathcal{Z}})\xrightarrow{x}H_{1}^{\mathrm{dR}}(\mathbf{G}|_{\mathcal{Z}}),$
which maps $\mathrm{Fil}(\mathbf{A}_{0}|_{\mathcal{Z}})$ to $\mathrm{Fil}(\mathbf{G}|_{\mathcal{Z}})$. By Grothendieck-Messing theory for $1$-motives, this morphism admits a canonical extension to a morphism between vector bundles over $\tilde{\mathcal{Z}}$,
@@
which determines a morphism (still denoted by $\tilde{x}$)
[…]
Proposition 4.8.
The morphism (4.2) can be extended to a morphism between line bundles:
(4.11) $\operatorname{Fil}\left(\mathbf{A}_{0}\right)|_{\tilde{\mathcal{Z}}}\stackrel{{\scriptstyle\tilde{x}}}{{\rightarrow}}L_{\mathbf{G}}|_{\tilde{\mathcal{Z}}}$.
@@
The proposition can be proved similarly as Proposition 4.4. ∎
Now we consider a geometric point $z$ on $\mathcal{B}_{\Phi}$. We can choose an étale neighborhood $X^{(z)}$ of $z$ in $\mathcal{C}^{*}_{\Phi}$ together with an automorphism $\gamma:\mathrm{Spec}\,\hat{R}_{z}\rightarrow\mathrm{Spec}\,\hat{R}_{z}$ where $\hat{R}_{z}$ is the completed local ring of $z$ in $\mathcal{C}^{*}_{\Phi}$ such that Proposition 2.9 holds. In the following discussion, the morphism $\mathrm{Spec}\,\hat{R}_{z}\rightarrow\mathcal{S}^{*}$ is meant to be the composition $\mathrm{Spec}\,\hat{R}_{z}\rightarrow X^{(z)}\rightarrow\mathcal{S}^{*}$ in the context of §2.7.
Proposition 4.9.
There is an isomorphism $\mathcal{Z}^{*}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}\xrightarrow{\sim}\mathcal{Z}^{*}_{\Phi}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}$ such that the following diagram is Cartesian.
(4.12) $\mathcal{Z}^{*}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}$$\mathrm{Spec}\,\hat{R}_{z}$$\mathcal{Z}^{*}_{\Phi}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}$$\mathrm{Spec}\,\hat{R}_{z}.$$\scriptstyle{\gamma}$
Proof.
Since both horizontal morphisms in (4.12) are finite and unramified, after passing to a further étale cover if necessary, we can assume the horizontal morphisms in (4.12) are closed immersion when restricted to each connected component, and the conclusion of Proposition 2.9 still holds.
Recall that $\gamma$ induces the identity on $\mathrm{Spec}\,\hat{R}_{z}/I_{z}$ (see Proposition 2.9) and $G^{\sharp}|_{\mathrm{Spec}\,\hat{R}_{z}/I_{z}}=\mathbf{G}_{\mathrm{Spec}\,\hat{R}_{z}/I_{z}}$ where $G^{\sharp}$ is the tautological semi-abelian scheme over $\mathcal{B}^{\Phi}$ and $\mathbf{G}$ is the universal semi-abelian scheme over $\mathcal{S}^{*}$. Hence
$\mathcal{Z}^{*}(m)|_{\mathrm{Spec}\,\hat{R}_{z}/I_{z}}=\mathcal{Z}^{*}_{\Phi}(m)|_{\mathrm{Spec}\,\hat{R}_{z}/I_{z}}$
by the definition of both sides. Suppose $x\in\operatorname{Hom}(\mathbf{A}_{0}|_{z},G^{\sharp}|_{z})$ such that $(x,x)=m$. Define a functor $\mathcal{Z}^{*}_{\Phi}(x)$ over $\mathrm{Spec}\,\hat{R}_{z}$ such that for any $\mathrm{Spec}\,\hat{R}_{z}$-scheme $S$, $\mathcal{Z}^{*}_{\Phi}(x)(S)$ is the isomorphism classes of tuples
$\left(\mathbf{A}_{0,S},\iota_{0,S},\lambda_{0,S},G_{S}^{\sharp},\iota_{S}^{\sharp},\lambda_{S}^{\sharp},\mathcal{F}_{S}^{\sharp},{\bf x}\right)$
where $(\mathbf{A}_{0,S},\ldots,\mathcal{F}_{S}^{\sharp})$ is the pullback of the tautological object over $\mathcal{C}^{*}_{\Phi}$ via $S\rightarrow\mathrm{Spec}\,\hat{R}_{z}\rightarrow\mathcal{C}^{*}_{\Phi}$ while ${\bf x}\in\operatorname{Hom}(\mathbf{A}_{0,S},G_{S}^{\sharp})$ and restricts to $x\in\operatorname{Hom}(\mathbf{A}_{0}|_{z},G^{\sharp}|_{z})$. By Grothendieck-Messing theory, $\mathcal{Z}^{*}_{\Phi}(x)$ is a subscheme of $\mathrm{Spec}\,\hat{R}_{z}$. By Lemma 4.6, we have the following decomposition of $\mathcal{Z}^{*}_{\Phi}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}$
$\mathcal{Z}^{*}_{\Phi}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}=\bigsqcup_{\begin{subarray}{c}x\in\operatorname{Hom}(\mathbf{A}_{0}|_{z},G^{\sharp}|_{z})\\ (x,x)=m\end{subarray}}\mathcal{Z}^{*}_{\Phi}(x).$
Similarly define a functor $\mathcal{Z}^{*}(x)$ over $\mathrm{Spec}\,\hat{R}_{z}$ such that for any $\mathrm{Spec}\,\hat{R}_{z}$-scheme $S$, $\mathcal{Z}^{*}(x)(S)$ is the isomorphism classes of tuples
$\left(\mathbf{A}_{0,S},\iota_{0,S},\lambda_{0,S},\mathbf{G}_{S},\iota_{S},\lambda_{S},\mathcal{F}_{S},\mathbf{y}\right)$
where $(\mathbf{A}_{0,S},\ldots,\mathcal{F}_{S})$ is the pullback of the universal object over $\mathcal{S}^{*}$ via $S\rightarrow\mathrm{Spec}\,\hat{R}_{z}\rightarrow\mathcal{S}^{*}$ while $\mathbf{y}\in\operatorname{Hom}(\mathbf{A}_{0,S},\mathbf{G}_{S})$ and restricts to $x\in\operatorname{Hom}(\mathbf{A}_{0}|_{z},G^{\sharp}|_{z})$. Then $\mathcal{Z}^{*}(x)$ is a subscheme of $\mathrm{Spec}\,\hat{R}_{z}$, and we have the following decomposition of $\mathcal{Z}^{*}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}$
$\mathcal{Z}^{*}(m)|_{\mathrm{Spec}\,\hat{R}_{z}}=\bigsqcup_{\begin{subarray}{c}x\in\operatorname{Hom}(\mathbf{A}_{0}|_{z},G^{\sharp}|_{z})\\ (x,x)=m\end{subarray}}\mathcal{Z}^{*}(x).$
Take $S=\mathcal{Z}^{*}_{\Phi}(x)$. Recall that by taking the quotient of $(G_{S}^{\sharp},\iota_{S}^{\sharp},\lambda_{S}^{\sharp},\mathcal{F}_{S}^{\sharp})$ by the image of the period map $\underline{\mathfrak{n}}\xrightarrow[]{u}G_{S}^{\sharp}$, we get a degenerating abelian scheme $({}^{\heartsuit}G_{S},{}^{\heartsuit}\iota_{S},{}^{\heartsuit}\lambda_{S},{}^{\heartsuit}\mathcal{F}_{S})$ of type $\Phi$ relative to $(S,S\cap\mathcal{B}_{\Phi},\eta_{S})$. Let $\mathbf{y}$ be the composition of $\bf x$} with quotient map $G_{S}^{\sharp}\rightarrow{}^{\heartsuit}G_{S}$, then $\mathbf{y}\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}\left(\mathbf{A}_{0,S},{}^{\heartsuit}G_{S}\right)$ and $(\mathbf{y},\mathbf{y})=m$. Combining with Proposition 2.9, we have a tuple $\gamma^{*}\left(\mathbf{A}_{0,S},\iota_{0,S},\lambda_{0,S},{}^{\heartsuit}G_{S},{}^{\heartsuit}\iota_{S},{}^{\heartsuit}\lambda_{S},{}^{\heartsuit}\mathcal{F}_{S},\mathbf{y}\right)$ over $\gamma^{*}(\mathcal{Z}_{\Phi}(x))$ which satisfies the definition of $\mathcal{Z}^{*}(x)$. This shows that $\gamma^{*}(\mathcal{Z}^{*}_{\Phi}(x))$ is a substack of $\mathcal{Z}^{*}(x)$.
Let $I$ (resp. $J$) be the defining ideal of $\mathcal{Z}^{*}(x)$ (resp. $\gamma^{*}(\mathcal{Z}_{\Phi}(x))$) in $\mathrm{Spec}\,\hat{R}_{z}$. Then $I\subseteq J$. Since $\dim G^{\sharp}_{S}=\dim{}^{\heartsuit}G_{S}$, we have a canonical isomorphism $\operatorname{Lie}G^{\sharp}_{S}\cong\operatorname{Lie}{}^{\heartsuit}G_{S}$. By Proposition 2.9, $\gamma^{*}(\operatorname{Lie}{}^{\heartsuit}G_{S})\cong\operatorname{Lie}\mathbf{G}_{S}$. Combining with Corollary 3.7 and (4.7), we have
$L_{\mathbf{G}}|_{\gamma^{*}(S)}=\bar{\epsilon}\operatorname{Lie}\mathbf{G}_{\gamma^{*}(S)}\cong\gamma^{*}(\bar{\epsilon}\operatorname{Lie}{}^{\heartsuit}G_{S})\cong\gamma^{*}(\bar{\epsilon}\operatorname{Lie}G^{\sharp}_{S})=\gamma^{*}(L_{G^{\sharp}}|_{S}).$
Moreover by the definition of $\mathbf{y}$, the following diagram commutes
$H_{1}^{\mathrm{dR}}({\bf A}_{0}|_{\gamma^{*}(S)})$$\gamma^{*}(\operatorname{Lie}G^{\sharp}_{S})$$\operatorname{Lie}\mathbf{G}_{\gamma^{*}(S)}.$$\scriptstyle{\gamma^{*}(\mathbf{y})}$$\scriptstyle{\gamma^{*}({\bf x})}\scriptstyle{\cong}$
By Proposition 4.4, the above diagram factors through
$H_{1}^{\mathrm{dR}}({\bf A}_{0}|_{\gamma^{*}(S)})$$\gamma^{*}(L_{G^{\sharp}}|_{S})$$L_{\mathbf{G}}|_{\gamma^{*}(S)}.$$\scriptstyle{\gamma^{*}(\mathbf{y})}$$\scriptstyle{\gamma^{*}({\bf x})}\scriptstyle{\cong}$
Let $\widetilde{\mathcal{Z}^{*}(x)}$ be the first order thickening of $\mathcal{Z}^{*}(x)$ in $\mathrm{Spec}\,\hat{R}_{z}$. By Grothendieck-Messing theory, the right-downward arrow in the above diagram extends over $\widetilde{\mathcal{Z}^{*}(x)}$, hence the whole diagram extends to
(4.13) ${H_{1}^{\mathrm{dR}}({\bf A}_{0}|_{\widetilde{\mathcal{Z}^{*}(x)}})}$$\gamma^{*}(L_{G^{\sharp}})|_{\widetilde{\mathcal{Z}^{*}(x)}}$$L_{\mathbf{G}}|_{\widetilde{\mathcal{Z}^{*}(x)}}$$\scriptstyle{\gamma^{*}(\mathbf{y})}$$\scriptstyle{\gamma^{*}({\bf x})}\scriptstyle{\cong}$
By Proposition 4.8 and Proposition 4.4, $\mathcal{Z}^{*}(x)$ (resp. $\mathcal{Z}^{*}_{\Phi}(x)$) is the vanishing locus of $\gamma^{*}(\mathbf{y})$ (resp. $\gamma^{*}({\bf x})$) in diagram (4.13). This shows that
$\widetilde{\mathcal{Z}^{*}(x)}\cap\mathcal{Z}^{*}_{\Phi}(x)=\mathcal{Z}^{*}(x),$
in other words, $J\equiv I\pmod{I^{2}}$. In particular $J\subseteq I$. Hence $I=J$, and $\gamma^{*}(\mathcal{Z}^{*}_{\Phi}(x))=\mathcal{Z}^{*}(x)$. The proposition is proved. ∎
5. Pullback of algebraic special divisors
5.1. Morphisms betweeen unitary Shimura varieties